# A kinematic origin of the white dwarfs in the Solar neighborhood

### v1

@@ 5.1 Orbital analysis of the Solar neighborhood
5.1.1 The average Solar neighborhood lifetime
In this first exercise, we have counted the stars that exit the Solar neighborhood as a function of time. The goal is to measure what is the rate of “renewal” of white dwarfs within this volume. Fig. 1 shows the results for our two reference samples of white dwarfs, the Gaia observed sample (dashed lines) and the synthetic one (continuous lines; see Sections 2 and 3, respectively), color-coded according to the different Galactic components considered.
The analysis of the results clearly reveals that the fraction of the initial number of stars (-axis of Fig. 1) rapidly decreases. In less than 10 Myr, more than 90% of the initial number of white dwarfs have left our Solar neighborhood. This fact can be quantified by assuming an exponential decay, and computing then the lifetime, i.e. the elapsed time for the star fraction to decrease by an factor. The results are given in Table 2 for the different populations considered. As expected, the dispersion effect is more pronounced for thick disk objects than thin disk ones, and even more for halo objects that possess, on average, larger velocity dispersion relative to the Sun. In particular, in approximately 3.3 Myr the majority of thin disk white dwarfs have left our Solar neighborhood, while the thick disk stars are doing it in Myr, and the halo stars scatter off in less than 1 Myr.
@@ 5.1.1 The average Solar neighborhood lifetime
We also stress that the exclusion/inclusion of low-speed objects ( km s-1; see Section 2) from the observed and simulated samples does not change the results by more than 1 %. Moreover, we note that the candidates in the observed sample have been classified using a Random Forest algorithm, which tends to under represent low-speed objects for the Galactic thick disk and halo components by classifying them as thin disk stars. Consequently, both factors tend to increase the average speed of the different Galactic components of the observed sample.
We conclude that the mean lifetime of the different white dwarf populations is very short compared to the orbital period of the Sun around the Galaxy. Thus, the concept of white dwarf Solar neighborhood should be understood as an ephemeral concept, as there are no physical processes in the lives of these stars, except for the planetary nebula phase, with a duration shorter than or of the order of a couple of million years. Nevertheless, taking into account the long evolutionary timescales of white dwarfs and assuming a space density that is constant at low Galactic latitudes, the population that is currently residing in the Solar neighborhood is assumed to be representative for statistical studies (see Section 5.2.1 and 5.2.3).
@@ 5.1.2 White dwarf birth places
Identifying the exact coordinates where individual stars come from is indeed a futile endeavor for objects as old as the white dwarfs. However, we can study the regions of the Milky Way where each population spends most of their time, and use this information as a proxy for their most likely birth places. In order to do this, we computed the orbits for nearly white dwarf for each Galactic component and adopted a 2 Gyr time interval to integrate the orbits of the synthetic sample. Such a long time interval (in comparison with the dissipative time previously estimated) ensures a full description of the parameter space that is covered by each star of our sample.
In Fig. 2 we show the spatial Galactocentric distributions covered by the orbits of the thin and thick disk, and halo populations. For each Galactic component we represent them in a corner plot (where the top panels correspond to histogram distributions, while the rest of the panels show the density plots of the different coordinate projections). As expected, halo members (bottom corner plot) show a wider spatial distribution within our Galaxy than the disk components, while thin disk (top corner plot) objects are concentrated in a more defined ring in the plane than thick disk population (middle corner plot).
A closer look at the top corner panel in Fig. 2 reveals the small deviation the thin disk members have with respect to an almost-circular orbit, which are due to their small velocity dispersion. The projection shows a truly thin ring, with a kpc width at the level, which is indicative of the fact that the orbits are mostly circular. There is also a narrow span in in which they merely deviate from the disk plane ( 330 pc also at the level), reaching at most 1 kpc in a few extreme cases.
With respect to the thick disk population (middle corner plot), the results reveals a wider spatial distribution of stars, reaching larger values in ( 780 pc at ). Whereas the maximum deviations from the Solar location in and are similar to those of the thin disk, the deviations in are almost the double. The “ring” in the projection of thick disk orbits broadens, reaching 3.4 kpc, suggesting more eccentric orbits associated to the members of the thick disk population.
@@
Complementary to the spatial distribution discussed, we computed the guiding-center radius, , defined as the radius that a circular orbit would have for a given component of the orbital angular momentum. The guiding-center radius assumes the validity of the epicyclic approximation for the stellar motion in the Galactic potential and its a good proxy for the mean orbital motion of a given star (Kordopatis et al., 2017). It is suggested that the effects of increased eccentricity and radial migrations due to interactions at the Lindblad resonances and the corotation radius in the Galactic disk (“blurring”, Sellwood & Binney, 2002; Schönrich & Binney, 2009) can be removed by investigating the properties of stars as a function of their guiding-center radii (Kordopatis et al., 2017). One way to calculate this value is by means of a galpy
function defined as, , where is the Galactocentric radius, and is the circular velocity at the Solar radius (i.e. the component in the Cartesian reference frame).
In Fig. 3 we represent the distribution of for the different Galactic components of our synthetic population. As a visual aid, we represent the position of the Sun (kpc) by a vertical red dashed line. Table 3 lists the mean values and standard deviations for these distributions.
The guiding-center radius distribution of the thin disk population is, as expected, centered at the Sun’s location with a relative small dispersion of values both towards and outwards the Galactic center. This result suggests that 68% of stars in the Solar neighborhood where formed at less than 1 kpc away from the Sun. We note a similarly symmetric distribution for the thick disk population, although centered at a smaller radius, implying that those objects with suffered with radial migration of their average orbits. Finally, the halo population shows a much wider distribution than the disk populations. Even more, halo stars spend longer time closer to the Galactic center than their current position, i.e., 92.16 % of them have a kpc. Thus, the halo population that we currently see in the Solar neighborhood belongs to what is likely the stellar components of the “inner halo”, given that their orbits are mostly planar and do not extend beyond –25 kpc. The large range of that the synthetic population present is due to the larger Galactic orbit eccentricities that are measured for the observed halo stars (e.g., Torres et al., 2019b; Raddi et al., 2022). Their immersion into the inner regions of the Milky Way is also indicative of stronger interactions with the Galactic bar and spiral arms, which may have scattered them from initially circular disk orbits.
@@ 5.1.3 Distribution of orbital parameters
Thin disk stars (top panel of Fig. 4), for which both synthetic and observed samples seem to be in a reasonable good agreement, present high values and small eccentricities, both features strongly related to circular orbits in the disk plane. Thick disk members (middle panel) slightly tend to more eccentric orbits. In this case, when comparing observed and synthetic samples, there is a clear lack of low eccentric thick disk orbits in the former one. As stated in Section 2, this fact is attributed to the higher probability of the Random Forest algorithm classifying low-speed circular orbits as thin disk stars. Finally, the distribution of halo objects (bottom panel) results in a broader range of angular momenta for the observed sample compared to the synthetic sample. There are even some observed halo members that have negative , indicating that they rotate in the opposite direction with respect to the Galactic rotation (retrograde orbits) and have very low angular momenta.
@@
We note an intuitively clear relation between high and low . Provided that a very elliptic orbit has a lower energy than the corresponding circular orbit, both crossing the Solar circle at 8 kpc, the elliptic orbit will have its apocenter in the Solar neighbourhood end, thus, its pericentre will be closer to the Galactic center. For this reason, there will be a large number of stars passing through inner regions. Simulated halo stars that display small orbit eccentricity have that are more concentrated towards the Solar circle, where circular orbits are usual. The correlation between and , bottom panel in Fig. 5, shows that all the retrograde orbits ( kpc km s-1) distinctly cluster at kpc. This result could imply that the few observed white dwarfs with retrograde orbits could have acquired their highly eccentric orbits after being scattered from the inner regions of the Milky Way. Nevertheless, also stellar streams that have formed via perturbations due to past encounters of the Milky Way with other satellite galaxies are characterized by retrograde orbits (Helmi et al., 2018; Myeong et al., 2019). On the other hand, halo stars with prograde orbits ( kpc km s-1) have a wider distribution of that mimics that of the overall subgroup, indicating a variety of possible origins, in which interactions due to the outer Lindblad resonance of the Galactic bar can play a major role. In this context, it is worth comparing the distribution of orbital energy and -component of angular momentum for the simulated and observed white dwarfs as shown in Fig. 6. While thin and thick disk members occupy a limited portion of the diagrams, which is related to their more circular orbits as it is shown in Fig. 4, the halo white dwarfs can display a much larger variety of angular momenta and energies. The latter group consists of 89 stars that were classified by Torres et al. (2021), with 11 of them possessing unusually young ages of less than 7 Gyr. This subgroup of stars appears to possess comparatively larger energies for a given value of , placing most of them beyond the 1 contour of Fig. 6. While there is an extremely very low chance of contamination with the most extreme thick disk stars, those authors suggested that these 11 white dwarfs could possess halo kinematics, despite their young age, due to more eventful lives as binary-star merger products or relics of past mergers of the Milky Way with its satellite galaxies.
@@ 5.2.1 The Solar neighborhood size
The exact size of the Solar neighborhood is an arbitrary value. It depends on which stellar objects we are considering and what the ultimate purpose of the study is. In the case of the white dwarf population, the concept of the Solar neighborhood can vary from a few tens to a few hundred parsecs. For instance, one can find studies in the literature on the star formation history up to 40 pc (Cukanovaite et al., 2023), the age-metallicity relation up to a few hundred pc (Rebassa-Mansergas et al., 2021), or the white dwarf spectral type-temperature distribution up to 500 pc (Torres et al., 2023).
@@
On the other hand, we have already checked in Section 5.1.1 that the Solar neighborhood is a transient concept. We now address the question of whether the different volumes that we may consider as the Solar neighborhood are equivalent, that is, if the objects that make up this sample are representative of the same region of the Galaxy.
@@
To analyze how the size of the Solar neighborhood affects the spatial distribution, we have simulated volumes of 40, 100, 500 pc and an extra large volume of 1.5 kpc, assuming in all cases the same velocity distribution as for the 100 pc sample. In Table 5 we summarized the two main parameters that characterize the spatial distribution of the white dwarf Galactic components: the width of the ring, , and the width of the perpendicular layer with respect to the Galactic plane, .
@@
The analysis of Table 5 reveals that for Solar neighborhood volumes up to 500 pc, differences in for the Galactic spatial distribution of white dwarfs are minor, while the experiences a 30–50 % increase for the thin and thick disk populations, respectively. This implies that current white dwarf surveys up to 500 pc (for instance, practically 90% of Gaia spectra are contained within that volume) are representative of the same putative Galactic white dwarf population. In other words, the Galactic origin of the vast majority of white dwarf within 40, 100 or 500 pc from the Sun can be assumed to be the same. Hence, the corresponding star formation histories, luminosity functions, mass distributions or other ensemble properties extracted from the different samples can be considered equivalent.
@@
However, that is not the case for the largest sample analyzed, the 1.5 kpc volume. In particular, concerning the disk populations, the Galactic region contributing to a 1.5 kpc Solar neighborhood radius is considerably larger than in previous cases. For instance, the ring width has increased by approximately a factor of 1.3 compared to the 100 pc sample, and the dispersion in coordinate nearly doubled. As a result, future analyses of this larger sample should consider incorporating a wider variety of stars, such as those with different metallicities coming from a broader region of the Galaxy.
@@
Finally, we note that the halo population does not present a large difference in for the different neighborhood sizes and it is only characterized by an 30 % increase in , implying that the velocity distribution of inner-halo white dwarfs detected within 100-pc could also be representative for the much larger volume.
@@
5.2.2 The effect of null radial velocity assumption
We conducted a simulation to assess the impact of assuming null radial velocities on our estimation of the spatial distribution of the Solar neighborhood population. In this simulation, we artificially inflate the mean values and dispersions by one-third compared to those adopted in Table LABEL:table:vel-gauss (see discussion in Section 2 and Rowell & Kilic, 2019). The results corresponding to the spatial distribution parameters (ring width, , and -dispersion, , along with the mean values and standard deviations for the guiding-center radius distributions of each Galactic component) are summarized in Table 6. Additionally we selected a sample of 20 469 FGK-type main-sequence and sub-giant stars within 100 pc from the Sun, which have reliable Gaia astrometry and radial velocities (with the first 13 parameter flags of the flags_gspspec
string set equal to zero; Recio-Blanco et al. (2023)). This sample of not-evolved stars occupies the same volume as that of the observed white dwarfs and, thus, it offers a useful comparison for the bias introduced by accounting or not accounting for radial velocities of stars. As for the local white dwarf sample, there are mostly thin disk objects, but also some thick disk and halo stars are present. As for the white dwarfs, we analyzed these stars by integrating their orbits but, in this case, also accounting for their measured radial velocities. The results are listed at the last row of Table 6.
From the comparison with the previously obtained results (see Table 3 and Table 5 for the 100 pc simulation), we deduced that inflating the space velocity by one-third increases the volume occupied by the orbits of thin disk white dwarfs by 30–35 %, as quantified by and , and reduces by 2% the mean with a 40 % larger . A similar result, but of the order of 10–15 %, is found when considering the main-sequence and sub-giant sample, which naturally incorporates measurements of the radial velocity. However, the inclusion of radial velocities would introduce a much more significant change in the volume occupied by thick-disk and halo white dwarfs, which appear to double in size for the latter sub-sample that is characterized by more extreme kinematics. The larger reduction of guiding radius of halo white dwarfs is, thus, accompanied by an increment of the eccentricity of their orbits. A consequence of these changes is % reduction of the Solar neighborhood lifetime with respect to the results of Section 5.1.1, corresponding to 2.47, 1.04, and 0.48 Myr for the thin and thick disk, and halo, respectively.
In summary, we conclude that the inclusion of an estimate of the radial component of the velocity of white dwarfs is currently rather uncertain but, if their kinematics is similar to that of main-sequence and sub-giants belonging to the same sub-populations, our results will be valid within a systematic 15–30 % uncertainty.

### v2

@@ 5.1 Orbital analysis of the solar neighborhood
5.1.1 Average solar neighborhood lifetime
In this first exercise, we counted the stars that exit the solar neighborhood as a function of time. The goal is to measure what is the rate of “renewal” of white dwarfs within this volume. Figure 1 shows the results for our two reference samples of white dwarfs, the Gaia observed sample (dashed lines) and the synthetic one (continuous lines; see Sections 2 and 3, respectively), color-coded according to the different Galactic components considered.
The analysis of the results clearly reveals that the fraction of the initial number of stars (-axis of Fig. 1) rapidly decreases. In less than 10 Myr, more than 90% of the initial number of white dwarfs have left our solar neighborhood. This fact can be quantified by assuming an exponential decay and computing then the lifetime, namely, the elapsed time for the star fraction to decrease by an factor. The results are given in Table 2 for the different populations considered. As expected, the dispersion effect is more pronounced for thick disk objects than thin-disk ones and even more for halo objects that possess, on average, larger velocity dispersion relative to the Sun. In particular, in approximately 3.3 Myr the majority of thin-disk white dwarfs have left our solar neighborhood, while the thick disk stars are doing it in Myr and the halo stars scatter off in less than 1 Myr.
@@ 5.1.1 Average solar neighborhood lifetime
We also stress that the exclusion or inclusion of low-speed objects ( km s-1; see Section 2) from the observed and simulated samples does not change the results by more than 1 %. Moreover, we note that the candidates in the observed sample have been classified using a RF algorithm, which tends to under represent low-speed objects for the Galactic thick disk and halo components by classifying them as thin-disk stars. Consequently, both factors tend to increase the average speed of the different Galactic components of the observed sample.
We conclude that the mean lifetime of the different white dwarf populations is very short compared to the orbital period of the Sun around the Galaxy. Thus, the concept of white dwarf solar neighborhood should be understood as an ephemeral concept, as there are no physical processes in the lives of these stars, except for the planetary nebula phase, with a duration shorter than or on the order of a couple of million years. Nevertheless, in taking into account the long evolutionary timescales of white dwarfs and assuming a space density that is constant at low Galactic latitudes, the population that is currently residing in the solar neighborhood is assumed to be representative for statistical studies (see Sects. 5.2.1 and 5.2.3).
@@ 5.1.2 White dwarf birth places
Identifying the exact coordinates where individual stars come from is indeed a futile endeavor for objects as old as the white dwarfs. However, we can study the regions of the Milky Way where each population primarily resides and use this information as a proxy for their most likely birth places. In order to do this, we computed the orbits for nearly white dwarf for each Galactic component and adopted a 2 Gyr time interval to integrate the orbits of the synthetic sample. Such a long time interval (in comparison with the dissipative time previously estimated) ensures a full description of the parameter space that is covered by each star of our sample.
In Fig. 2, we show the spatial Galactocentric distributions covered by the orbits of the thin and thick disk, and halo populations. For each Galactic component, we represent them in a corner plot (where the top panels correspond to histogram distributions, while the rest of the panels show the density plots of the different coordinate projections). As expected, halo members (bottom corner plot) show a wider spatial distribution within our Galaxy than the disk components, while thin-disk (top corner plot) objects are concentrated in a more defined ring in the plane, as compared to the thick-disk population (middle corner plot).
A closer look at the top corner panel in Fig. 2 reveals the small deviation the thin-disk members have with respect to an almost-circular orbit, which are due to their small velocity dispersion. The projection shows a truly thin ring, with a kpc width at the level, which is indicative of the fact that the orbits are mostly circular. There is also a narrow span in in which they merely deviate from the disk plane ( 330 pc also at the level), reaching 1 kpc at most in a few extreme cases.
With respect to the thick-disk population (middle corner plot), the results reveals a wider spatial distribution of stars, reaching larger values in ( 780 pc at ). Meanwhile, the maximum deviations from the solar location in and are similar to those of the thin disk, the deviations in are almost the double. The “ring” in the projection of thick disk orbits broadens, reaching 3.4 kpc, suggesting more eccentric orbits associated to the members of the thick disk population.
@@
As a complement to the spatial distribution discussed, we computed the guiding-center radius, , defined as the radius that a circular orbit would have for a given component of the orbital angular momentum. The guiding-center radius assumes the validity of the epicyclic approximation for the stellar motion in the Galactic potential and its a good proxy for the mean orbital motion of a given star (Kordopatis et al., 2017). It is suggested that the effects of increased eccentricity and radial migrations due to interactions at the Lindblad resonances and the corotation radius in the Galactic disk (“blurring;” Sellwood & Binney, 2002; Schönrich & Binney, 2009) can be removed by investigating the properties of stars as a function of their guiding-center radii (Kordopatis et al., 2017). One way to calculate this value is by means of a galpy
function defined as , where is the Galactocentric radius, and is the circular velocity at the solar radius (i.e., the component in the Cartesian reference frame).
In Fig. 3, we represent the distribution of for the different Galactic components of our synthetic population. As a visual aid, we represent the position of the Sun (kpc) by a vertical red dashed line. Table 3 lists the mean values and standard deviations for these distributions.
The guiding-center radius distribution of the thin-disk population is, as expected, centered at the Sun’s location with a relative small dispersion of values both towards and outwards the Galactic center. This result suggests that 68% of stars in the solar neighborhood where formed at less than 1 kpc away from the Sun. We note a similarly symmetric distribution for the thick disk population, although centered at a smaller radius, implying that those objects with suffered with radial migration of their average orbits. Finally, the halo population shows a much wider distribution than the disk populations. Even more, halo stars spend longer time closer to the Galactic center than their current position, namely, 92.16 % of them have a kpc. Thus, the halo population that we currently see in the solar neighborhood belongs to what is likely the stellar components of the “inner halo,” given that their orbits are mostly planar and do not extend beyond –25 kpc. The broad range of that the synthetic population present is due to the larger Galactic orbit eccentricities that are measured for the observed halo stars (e.g., Torres et al., 2019b; Raddi et al., 2022). Their immersion into the inner regions of the Milky Way is also indicative of stronger interactions with the Galactic bar and spiral arms, which may have scattered them from initially circular disk orbits.
@@ 5.1.3 Distribution of orbital parameters
Thin-disk stars (top panel of Fig. 4), for which both synthetic and observed samples appear to be in a reasonable good agreement, present high values and small eccentricities. Both features are strongly related to circular orbits in the disk plane. Thick disk members (middle panel) slightly tend to more eccentric orbits. In this case, when comparing observed and synthetic samples, there is a clear lack of low eccentric thick disk orbits in the former one. As stated in Section 2, this fact is attributed to the higher probability of the RF algorithm classifying low-speed circular orbits as thin-disk stars. Finally, the distribution of halo objects (bottom panel) results in a broader range of angular momenta for the observed sample compared to the synthetic sample. There are even some observed halo members that exhibit a negative , indicating that they rotate in the opposite direction with respect to the Galactic rotation (retrograde orbits), with very low angular momenta.
@@
We note an intuitively clear relation between high and low . Provided that a very elliptic orbit has a lower energy than the corresponding circular orbit, both crossing the solar circle at 8 kpc, the elliptic orbit will have its apocenter in the solar neighbourhood end; thus, its pericenter will be closer to the Galactic center. For this reason, there will be a large number of stars passing through inner regions. Simulated halo stars that display small orbit eccentricity with values that are more concentrated towards the solar circle, where circular orbits are usual. The correlation between and , bottom panel in Fig. 5, shows that all the retrograde orbits ( kpc km s-1) distinctly cluster at kpc. This result could imply that the few observed white dwarfs with retrograde orbits could have acquired their highly eccentric orbits after being scattered from the inner regions of the Milky Way. Nevertheless, also stellar streams that have formed via perturbations due to past encounters of the Milky Way with other satellite galaxies are characterized by retrograde orbits (Helmi et al., 2018; Myeong et al., 2019). On the other hand, halo stars with prograde orbits ( kpc km s-1) have a wider distribution of that mimics that of the overall subgroup, indicating a variety of possible origins; these would involve interactions attributed to the outer Lindblad resonance of the Galactic bar playing a major role. In this context, it is worth comparing the distribution of orbital energy and -component of angular momentum for the simulated and observed white dwarfs as shown in Fig. 6. While thin and thick disk members occupy a limited portion of the diagrams, which is related to their more circular orbits as it is shown in Fig. 4, the halo white dwarfs can display a much larger variety of angular momenta and energies. The latter group consists of 89 stars that were classified by Torres et al. (2021), with 11 of them possessing unusually young ages of less than 7 Gyr. This subgroup of stars appears to possess comparatively larger energies for a given value of , placing most of them beyond the 1 contour of Fig. 6. While there is an extremely very low chance of contamination with the most extreme thick disk stars, those authors suggested that these 11 white dwarfs could possess halo kinematics, despite their young age, due to more eventful lives as binary-star merger products or relics of past mergers of the Milky Way with its satellite galaxies.
@@ 5.2.1 The solar neighborhood size
The exact size of the solar neighborhood is an arbitrary value. It depends on which stellar objects we are considering and what the ultimate purpose of the study is. In the case of the white dwarf population, the concept of the solar neighborhood can vary from a few tens to a few hundred parsecs. For instance, there are studies available in the literature on the star formation history up to 40 pc (Cukanovaite et al., 2023), the age-metallicity relation up to a few hundred pc (Rebassa-Mansergas et al., 2021), and the white dwarf spectral type-temperature distribution up to 500 pc (Torres et al., 2023).
@@
On the other hand, as we discuss in Section 5.1.1, the solar neighborhood is a transient concept. We now address the question of whether the different volumes that we may consider as the solar neighborhood are equivalent, that is, if the objects that make up this sample are representative of the same region of the Galaxy.
@@
To analyze how the size of the solar neighborhood affects the spatial distribution, we simulated volumes of 40, 100, and 500 pc and an extra large volume of 1.5 kpc, assuming in all cases the same velocity distribution as for the 100 pc sample. In Table 5, we summarized the two main parameters that characterize the spatial distribution of the white dwarf Galactic components: the width of the ring, , and the width of the perpendicular layer with respect to the Galactic plane, .
@@
The analysis in Table 5 reveals that for solar neighborhood volumes up to 500 pc, differences in for the Galactic spatial distribution of white dwarfs are minor, while the values see a 30–50 % increase for the thin- and thick-disk populations, respectively. This implies that current white dwarf surveys up to 500 pc (e.g., practically 90% of Gaia spectra are contained within that volume) are representative of the same putative Galactic white dwarf population. In other words, the Galactic origin of the vast majority of white dwarf within 40, 100, or 500 pc from the Sun can be assumed to be the same. Hence, the corresponding star formation histories, luminosity functions, mass distributions, or other ensemble properties extracted from the different samples can be considered to be equivalent.
@@
However, that is not the case for the largest sample analyzed, the 1.5 kpc volume. In particular, with respect to the disk populations, the Galactic region contributing to a 1.5 kpc solar neighborhood radius is considerably larger than in previous cases. For instance, the ring width has an increased by approximately a factor of 1.3 compared to the 100 pc sample, while the dispersion in the coordinate has nearly doubled. As a result, future analyses of this larger sample should consider incorporating a wider variety of stars, such as those with different metallicities coming from a broader region of the Galaxy.
@@
Finally, we note that the halo population does not present a substantial difference in for the different neighborhood sizes and it is only characterized by an 30 % increase in , implying that the velocity distribution of inner-halo white dwarfs detected within 100-pc could also be representative for the much larger volume.
@@
Note: Adopted values correspond to the level. |
5.2.2 Effect of null radial velocity assumption
We conducted a simulation to assess the impact of assuming null radial velocities on our estimation of the spatial distribution of the solar neighborhood population. In this simulation, we artificially inflate the mean values and dispersions by one-third compared to those adopted in Table LABEL:table:vel-gauss (see discussion in Sects. 2 and Rowell & Kilic, 2019). The results corresponding to the spatial distribution parameters (ring width, , and -dispersion, , along with the mean values and standard deviations for the guiding-center radius distributions of each Galactic component) are summarized in Table 6. Additionally we selected a sample of 20 469 FGK-type main-sequence and sub-giant stars within 100 pc from the Sun, which have reliable Gaia astrometry and radial velocities (with the first 13 parameter flags of the flags_gspspec
string set equal to zero; Recio-Blanco et al. (2023)). This sample of non-evolved stars occupies the same volume as that of the observed white dwarfs and, thus, it offers a useful comparison for the bias introduced by accounting or not accounting for radial velocities of stars. As for the local white dwarf sample, it is mostly comprised of thin-disk objects, but also some thick-disk and halo stars are present. As for the white dwarfs, we analyzed these stars by integrating their orbits but, in this case, we also accounted for their measured radial velocities. The results are listed at the last row of Table 6.
From the comparison with the previously obtained results (see Tables 3 and 5 for the 100 pc simulation), we deduced that inflating the space velocity by one-third increases the volume occupied by the orbits of thin-disk white dwarfs by 30–35 %, as quantified by and , and reduces by 2% the mean with a 40 % larger . A similar result, but on the order of 10–15 %, is found when considering the main sequence and sub-giant sample, which naturally incorporates measurements of the radial velocity. However, the inclusion of radial velocities would introduce a much more significant change in the volume occupied by thick-disk and halo white dwarfs, which appear to double in size for the latter sub-sample that is characterized by more extreme kinematics. The greater reduction in the guiding radius of a halo white dwarf is thus accompanied by an increment of the eccentricity of their orbits. A consequence of these changes, there is a % reduction of the solar neighborhood lifetime with respect to the results of Section 5.1.1, corresponding to 2.47, 1.04, and 0.48 Myr for the thin-disk, thick-disk, and the halo, respectively.
In summary, we conclude that the inclusion of an estimate of the radial component of the velocity of white dwarfs is currently rather uncertain. However, if their kinematics is similar to that of main sequence and sub-giants belonging to the same sub-populations, our results will be valid within a systematic 15–30 % uncertainty.